OpenStudy (anonymous):

AB is tangent to circle O at B. If AB=9 and AO=21.6, what is the length of the radius (r)?

OpenStudy (jdoe0001):

got picture?

OpenStudy (anonymous):

Yeah @jdoe0001

OpenStudy (anonymous):

use Pythagoras's theorem

OpenStudy (anonymous):

I don't know how to do that very well and I really need help. @surjithayer @jdoe0001

OpenStudy (blank ):

Radius is 20

OpenStudy (jdoe0001):

 \(\bf c^2=a^2+b^2\implies \sqrt{c^2-b^2}=b\implies \sqrt{21.6^2-9^2}=a\) 

OpenStudy (blank ):

Thats is assuming that Angle ABO is a right angle

OpenStudy (jdoe0001):

it's since the "r" is a radius touching a tangent at the tangency point

OpenStudy (anonymous):

Radius is 20...?

OpenStudy (anonymous):

19.6? @jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thank you so much!! Could you help me with some others?